11 people plan a trip and hire a minibus that exactly accommodate all of them. If only 5 of the eleven people can drive, in how many ways can they be seated?
Answer by Edwin McCravy(20054)

Choose the driver any of (5C1) = 5 ways.

For each of the 5 choices of the driver, there are 10P10 
or 10! ways to seat the 10 passengers.

Answer (5C1)(10P10) = 5*10! = 5*3628800 = 1,8144,000 ways.

Edwin